OpenStudy (jamaljunior):

Two trains leave the same station at the same time, but are headed in opposite directions. One train travels at 70 mph and the other train travels at 80 mph. How much times passes until the trains are 600 mi apart?

OpenStudy (awesomedude2078):

@Awolflover1

OpenStudy (sshayer):

relative speed v=70+80=150mph distance=600 miles time=600/150=

OpenStudy (jamaljunior):

4

OpenStudy (sshayer):

4 hours is correct.

OpenStudy (jamaljunior):

thanks fm

OpenStudy (sshayer):

yw